OpenStudy (anonymous):

I need help finding the limit of t-t(sqrt(t))/2t^3/2 + 3t - 5 as x approaches infinity. I think I'm doing something wrong algebraically. The answer I came up with is 1/2. I think this answer is wrong. Please help by breaking down the steps so I can check my work. Thank you. :)

OpenStudy (anonymous):

Write this out properly using the equation tab or the draw tab. I really dont understand the function you have up there

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} (t-\sqrt{t})/2t ^{3/2}+3t-5\]\[t-\sqrt{t}\]

OpenStudy (anonymous):

Ok I tried using the equation box...I somehow accidentally inserted an extra t- square root of t below the actual equation...please ignore that part :)

OpenStudy (anonymous):

\[\lim_{t\to \infty}\frac{t-t\sqrt{t}}{2t^{\frac{3}{2}}-3t+5}\]

OpenStudy (anonymous):

that is my guess, and you are close, but i think you should get \(-\frac{1}{2}\)

OpenStudy (anonymous):

ok...thanks for the valuable input....I thank you...where do you think I missed the negative...I'm spacing out...I can't find it...?

OpenStudy (anonymous):

it is in front of the \(t\sqrt{t}\) term

OpenStudy (anonymous):

\[\frac{-t\sqrt{t}}{2t^{\frac{3}{2}}}=-\frac{1}{2}\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw